
- Found the OPTGUI reports TELESCOPE is down. OPTGUI reboot reports: 139.229.12.8: RPC: remote system error - No route to host - error: no response from TCS. J. Rojas found the support power supply is broken! He did tests of consumption, the Heurikon power supply needs an extra, external power supply and finally he added a new power supply. Then we recovered the OPTGUI and then the pointing.
- 08:00 UT OPTGUI START widget did not work, rebooted CHIRON.
| Total Times | Time Observed | Time Engineering | Time lost technical | Time lost weather | Time lost other | Total Program Time |
|---|---|---|---|---|---|---|
| 2021-01-18 | 6.5 | 0.0 | 2.0 | 0.0 | 0.0 | 8.5 |
Links
[1] mailto:thenry@astro.gsu.edu
[2] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/33336/delete?destination=print/content/enr-ctio-15-m-2021-01-18
[3] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/33336/edit?destination=print/content/enr-ctio-15-m-2021-01-18
[4] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/add/node/17438?destination=print/content/enr-ctio-15-m-2021-01-18
- J. Rojas requested: if V is < 4.8[v], we must report!
- The air conditioning unit was NOT working fine. T in range [10,35] C. Please, check it.
- 02:47 UT A small earthquake, impacts pointing on GJ205, second integration (file 1117.) Also, it looks like fiber center is not (291, 249) but (292, 249.)
- File 1125, telescope moved ~ 10" S! Added one integration to 'compensate'.
- The anti static wrist wrap is broken. A replacement would be appreciated.
- Did not observe NOS objects: 1,3,5,7,10,13 and ended obs on HR 5049 = 67 MUS at 09:25 UT with Sun at -7.6 degrees of elevation.